Jackson Automotive has net working capital of $22,600, current assets of $56,500, equity of $62,700, and long-term debt of $31,900. What is the amount of the net fixed assets?

Multiple Choice

  • $9,300

  • $49,400

  • $94,600

  • $103,900

  • $72,000
Answer #1

Answer

--Correct Answer = Option #4: Net Fixed assets = $ 72,000

--Calculation:
>Current Liabilities = Current asset - working capital = 56500 - 22600 = $ 33,900
>Total Liabilities = Long term debt + current liabilities = 31900 + 33900 = $ 65,800
>Total Assets = Total Liabilities + total equity = 65800 + 62700 = $ 128,500
>Total Fixed assets = Total Assets - Current Assets = $ 128500 - 56500 = $ 72,000 Answer